# Reinforcing Infrastructure Networks with Multicriteria Portfolio Decision Analysis: An Application to Railway Stations in Finland

### v1

@@ Abstract
Advanced societies are crucially dependent on critical infrastructure networks for the reliable delivery of essential goods and services. Hence, well-founded analyses concerning disruptions are needed to guide decisions that seek to ensure the performance of these networks in the face of failures caused by vulnerabilities to external hazards or technical malfunctions. In this setting, we develop a multicriteria decision analysis approach to support the formulation of cost-efficient portfolios of preventive reinforcement actions. Our approach is general in that it (i) allows for multiple objectives, such as those that represent the volume of traffic that is enabled between alternative origin-destination pairs in a transportation network, (ii) uses methods of probabilistic risk assessment to quantify the expected performance of the network, and (iii) solves optimization problems to identify those combinations of reinforcement actions that are cost-efficient in improving the performance of the network, given the available, possibly incomplete information about the relative importance of objectives. Our methodological contributions are illustrated by a case study on the analysis of railway switches at a representative Finnish railway station.
@@
Topological metrics can miss alignment with actual network performance goals.
Multicriteria analysis aids efficient resource allocation for risk management.
Algorithms proposed to identify cost-efficient portfolios.
Illustrative case study of railway stations in Finland.
@@ 1 Introduction
In Europe, rail networks amounted to 202,000 km in 2022, with considerable recent growth in high-speed trains. Disruptions in rail networks can undermine the achievement of overarching performance objectives, such as ensuring connectivity between strategic points or providing reliable connections for passengers and goods delivery. These disruptions can be caused by failures associated with the usual wear of technical systems or by vulnerabilities to external hazards, including extreme weather conditions and intentional attacks. As such, it is necessary to understand what types of disruption can affect the rail network, what impacts these disruptions can cause, and what actions are needed to mitigate them subject to the requirement that these actions be selected as cost-effectively as possible [see, e.g., 3, 4, 5].
@@
Not all components in transportation networks are equally important, as disruptions in some parts of the network will cause its performance to degrade more than those in others. Moreover, because impacts on network performance depend on the states of all network components, there is a need to evaluate the impacts of combinations of disruptions (e.g., a component may not be so important on its own, but its disruption simultaneously with another component may have drastic consequences for the performance of the network). Furthermore, it is pertinent to examine how the expected impacts of disruptions depend on alternative assumptions about the underlying probabilities of component failures, given that the information about the probabilities concerning component disruptions may be incomplete.
@@
Reinforcement actions reduce the possibility of disruptions that would undermine the performance of the network. Because many reinforcement actions are typically implemented jointly, we analyze portfolios consisting of many actions. These analyses support the cost-efficient allocation of available resources to select those portfolios of reinforcement actions, which, taken together, provide the highest expected performance relative to the costs of implementing these actions [6]. Furthermore, such analyses can be produced for a range of budget levels for aggregate implementation costs, for instance, to prepare plans for the cost-efficient allocation of additional resources or, in case of budget cuts, to identify which one of the previously planned reinforcement actions can be left out so that the impact on expected network performance will remain as small as possible.
@@
In this paper, we develop an approach based on multicriteria portfolio decision analysis in order (i) to assess and aggregate multiple objectives that reflect the services required of infrastructure networks and (ii) to guide the allocation of resources to actions that are cost-efficient in mitigating disruptions in different parts of the network. Specifically, we represent the network by nodes (components subject to failure) and edges (other components), and our formulation can be used to identify which nodes (or combinations of nodes) are most important to the network performance. Moreover, it helps identify which portfolios of reinforcement actions are cost-efficient in ensuring a desired level of network performance. Examples of such reinforcement actions include fortifying selected nodes against disruptions.
@@
The remainder of this paper is organized as follows. Section 2 discusses earlier approaches to analyzing disruptions and their impacts on the performance of transportation networks. Section 3 develops an approach that combines probabilistic risk assessment with multicriteria portfolio decision analysis to quantify the expected performance of the network with regard to several objectives and, in addition, guides the allocation of resources to those reinforcement actions that fortify them cost-efficiently. Section 4 presents a case study on the fortification of switches to improve the reliability of connections at a representative railway station in Finland. Section 5 discusses the numerical results, and Section 6 outlines future research areas.
@@ 2 Background
Early studies of railway networks focused mainly on their topological configuration due to limited data and computational power [7]. Later, there was a proliferation of specific models that have been developed to reduce travel times [8], design schedules [9], plan new lines [10], reduce operating costs [11], and secure the functionality of the tracks, among others. In many countries, infrastructure decisions (e.g., ensuring that the railway network is safe and functional) and operational decisions (e.g., ensuring that train timetables are being maintained) are made by entirely different entities. This is reflected in the fact that infrastructure and operational research have historically been somewhat decoupled. Within these streams of research, we focus on analyzing the network as a critical infrastructure in support of services.
@@
Most railway system analysis relies on network theory to identify critical components, evaluate the network performance, and develop strategies to reinforce or disrupt them [see, e.g., 12, 13, 14, 15]. Latora and Marchiori [16] present a method to find the critical components of an infrastructure network represented by nodes and edges. They also analyze how improvements, such as adding edges between nodes, enhance network performance. In their model, performance is measured using a topological metric, i.e., a metric that relies purely on how nodes are positioned and connected in the network, to quantify how efficiently information can be exchanged over the network. Ip and Wang [17] propose a methodology to evaluate the resilience of railway networks based solely on topological metrics, such as the number of independent paths. They also propose an algorithm to strengthen the network for greater resilience. Fecarotti et al. [18] propose a non-linear integer programming model that considers topological metrics to select maintenance strategies.
@@
Although several studies have used topological metrics to assess network performance, there has been less work on evaluating the quality of these assessments [19]. In this regard, a general drawback of topological metrics is that these metrics are not necessarily closely linked with the performance objectives of the network. For example, Hao et al. [20] propose a multi-objective optimization approach to identify critical components on a network, accounting for their interactions with other systems. One of their key findings is that the criticality of the nodes is not necessarily related to their topological importance. LaRocca et al. [21] conclude that topological metrics are of limited value in analyzing the robustness of power systems for different scenarios. Moreover, Alderson et al. [22] show that the criticality of a given node depends more on the set of disrupted nodes than on the node itself.
@@
Another challenge in evaluating the performance of networks under uncertainty is identifying possible hazards and their impacts. In some cases, it is unclear what hazards can affect the network [23] or to what extent the infrastructure will be capable of withstanding expected / unexpected hazards [24]. Zhang et al. [25] present a summary of recent studies to quantify the loss of functionality of railway systems due to various hazards. For combined hazards, they propose a framework to simulate functionality loss in a coupled railway and airline system. One of the key aspects in quantifying this loss is to account for the propagation of the disruption. In general, the assumption of independent failure rates leads to overestimating the network reliability, leading to poor design and unacceptable performance. Nazarizadeh et al. [26] propose a model which includes common cause failures and interactive failures. They apply the model to the Iranian railway system and show that it can provide better reliability estimates than other methodologies.
Several researchers have proposed approaches that use probabilities to quantify hazard scenarios. For example, Joshi et al. [27] and Yang et al. [28] consider scenarios of rainfall and tornadoes to assess the risks affecting the railway systems in India and China, respectively. Turoff et al. [29] propose a collaborative dynamic scenario model in which cascading effects due to interactions between critical infrastructure are estimated based on expert opinions. Zio [24] and Sedghi et al. [30] argue that there is a demand for analytical frameworks that help railway infrastructure managers understand and quantify the complexity of the network and, by doing so, help them prepare for hazards to ensure acceptable network performance. In this context, we develop a general assessment framework and accompanying algorithms that help assess the importance of components in a rail network and help to reinforce the network, of which they are an integral part.
3 Proposed Approach for Reinforcing Networks
@@ 3.1 Network Representation
Let $G(V;E)$ denote a network consisting of a set of nodes $V=\{1,\dots,n\}$ and a set of undirected edges $E\subseteq\{(i,i^{\prime})\mid i,i^{\prime}\in V\}$ between the nodes. A path is a sequence of nodes and edges that connect two nodes in the network. The state $x_{k}$ of node $k$ is a realization of a binary random variable $X_{k}$ such that $x_{k}=1$ if node $k\in V$ is disrupted and $x_{k}=0$ if it is operational. The state of the network is a vector that contains the states of all nodes $x=(x_{1},\dots,x_{n})\in\mathcal{X}=\{0,1\}^{n}$. Thus, the number of network states is $2^{n}$. If the disruptions of nodes occur independently, the uncertainties about the network state are fully characterized by the vector $p=(p_{1},\ldots,p_{n})$, where $p_{k}=\mathbb{P}[X_{k}=1]$ is the probability that node $k$ is disrupted. We consider that if a node is disrupted, none of the paths containing the node can be traversed. Thus, if the nodes in $D\subseteq V$ are disrupted, the remaining network is $G(V^{D};E^{D})$, where $V^{D}=V\setminus D$ and $E^{D}=\{(i,i^{\prime})\in E\mid i,i^{\prime}\in V^{D}\}\subseteq E$.
@@ 3.2 Assessing Network Performance
The performance of the network depends on its state. In many cases, this performance is multi-dimensional in that there are several objectives. For example, the performance of a transportation network can be assessed with regard to its ability to provide connections between relevant origin-destination pairs that are connected by the network. The objectives can also be based on topological metrics such as the average of the shortest distances between the nodes of the network [see, e.g., 31, 32].
@@
The attainment of multiple objectives can be measured by defining corresponding criteria and, more technically, by defining utility functions $u_{j}(\cdot):X\mapsto[0,1]$ such that $u_{j}(x)$ represents the utility that is associated with the performance on criterion $j=1,\dots,m$ when the network is in state $x\in\mathcal{X}$. We consider that these utilities cannot increase with the disruption of a component and normalize them so that the minimum $u_{j}(x)=0$ is associated with the states that give the worst performance on the $j$-th criterion (that is when all nodes are disrupted $x^{\circ}_{k}=1,k=1,\dots,n$). In contrast, the maximum $u_{j}(x)=1$ is attained when the network performs at its best on the $j$-th criterion (that is, when all nodes are operational $x^{*}_{k}=0,k=1,\dots,n$).
If the objectives are mutually preferentially independent (i.e., the preferences over a certain level of the objectives do not depend on the level of the other objectives as long as they stay the same), and single attributes are additive independent [see, e.g., 33], the preferences over the network performance associated with the network state $x\in\mathcal{X}$ can be represented by the additive utility function (1). However, because objectives are not necessarily equally important, we employ criterion weights that reflect the relative importance of these criteria, reflecting the overall utility increase when the criterion-specific performance changes from their minimum to the maximum level. The relative weight of the criterion $j=1,\dots,m$ is denoted by $w_{j}\in[0,1]$. Following the usual convention, we assume that these weights are normalized so that they sum to one. In the utility function (1), the additional utility brought about by an incremental improvement in the performance level of one criterion does not depend on the performance of the network on the other criteria.
$u(x,w)=\sum_{j=1}^{m}w_{j}u_{j}(x)\in[0,1].$ (1)
@@
However, obtaining a complete characterization of criteria weights can be challenging. For example, if the different criteria represent the planned volume of traffic between different origin-destination pairs in the network, it is possible that this volume is not fully known at the time of planning the fortification actions. In other cases, the decision-makers (DMs) may be reluctant to specify the importance of the criteria, or there can be situations with multiple DMs with different priorities. In response to this recognition, and also to support extensive sensitivity analyses, we characterize information about the relative importance of criteria through an information set $\mathcal{S}$ consisting of weight vectors that are compatible with the elicited preference statements [34]. These statements can be elicited by asking the DM to express statements concerning the relative importance of criteria. Thus, for example, if criterion 1 is at least as important as criterion 2 but not more than two times more important, the constraints $w_{2}\leq w_{1}\leq 2w_{2}$ hold. Assuming that the preferences are elicited through statements that correspond to linear constraints, then the resulting weight set $\mathcal{S}$ is a subset of the non-informative weight set $\mathcal{S}^{0}$
@@
3.3 Fortifying the Networks
The expected performance of the network can be improved by implementing reinforcement actions that, for example, lead to lower disruption probabilities at the nodes where they are implemented. Figure 1 illustrates the decision to implement an action to reinforce a network node. Without this action, the probability of disruption of node $k$ is $p_{k}$. If the corresponding action is implemented, this probability is reduced to $p^{\prime}_{k}<p_{k}$.
Figure 1: Decision tree for implementing an action to reinforce the network node $k$.
Usually, many actions can be carried out to improve performance. As a result, the probability of different levels of network performance depends on what is the portfolio (i.e., subset) of actions that are implemented. We consider $h$ possible binary reinforcement actions $q_{1},\ldots,q_{l}$ such that $q_{l}=1$ if action $l$ is implemented and $q_{l}=0$ if not, $l=1,\dots,h$. The portfolio of reinforcement actions is given by a vector $q=(q_{1},\dots,q_{h})\in\mathcal{Q}=\{0,1\}^{h}$. The total cost of a portfolio $q$ is $c_{tot}(q)\in\mathbb{R}_{+}$. In what follows, this total cost $c_{tot}$ is taken to be the sum of the costs of the individual actions $c_{l}$ in the portfolio, so that $c_{tot}(q)=\sum_{l=1}^{h}c_{l}q_{l}$. More complex cost structures can be easily introduced to account for cost synergies among actions.
A portfolio of reinforcement actions $q$ is feasible if (i) its total cost $c_{tot}(q)$ does not exceed the level of available resources $r$ (i.e., $c_{tot}(q)\leq r$) and (ii) satisfies relevant logical constraints (e.g., if actions $1$ and $2$ are mutually exclusive, allowing at most one of them to be implemented, the constraint $q_{1}+q_{2}\leq 1$ holds). The set of feasible portfolios is denoted by $\mathcal{Q}_{F}\subseteq\mathcal{Q}$. The analysis aims to determine those sets of feasible portfolios that satisfy the relevant constraints and contribute the most to attaining objectives, such as maximizing expected performance.

### v2

# Fortifying Critical Infrastructure Networks with Multicriteria Portfolio Decision Analysis: An Application to Railway Stations in Finland

@@ Abstract
Advanced societies are crucially dependent on critical infrastructure networks for the reliable delivery of essential goods and services. Hence, well-founded analyses concerning disruptions are needed to guide decisions that seek to ensure the performance of these networks in the face of failures caused by vulnerabilities to external hazards or technical malfunctions. In this setting, we develop an approach based on multicriteria decision analysis to support the identification of cost-efficient portfolios of preventive fortification actions. Our approach (i) allows for multiple objectives, such as those that represent the traffic volume that is enabled between alternative origin-destination pairs in a transportation network, (ii) uses methods of probabilistic risk assessment to quantify the expected performance of the network, and (iii) uses a search algorithm combined with an optimization model to identify those combinations of fortification actions that are cost-efficient in improving the performance of the network, given the available, possibly incomplete information about the relative importance of objectives and minimum performance requirements on them. Our methodological contributions are illustrated by a case study on the analysis of railway switches at a representative Finnish railway station.
@@
Topological metrics can miss alignment with network performance objectives.
Multicriteria analysis supports efficient resource allocation for risk management.
Search algorithms proposed to identify cost-efficient portfolios.
Illustrative case study on fortifying switches at railway stations in Finland.
@@ 1 Introduction
In Europe, rail networks amounted to 202,000 km in 2022, with considerable recent growth in high-speed trains [3]. Disruptions in rail networks can undermine related performance objectives, such as ensuring connectivity between strategic points or providing reliable connections for passengers and goods delivery. These disruptions can be caused by failures due to the usual wear of technical systems or by vulnerabilities to external hazards, including extreme weather conditions and intentional attacks. Consequently, there is a need to understand what types of disruption can affect the rail network, what impacts these disruptions can cause, and what actions are needed to mitigate them subject to the requirement that these actions must be selected as cost-effectively as possible given resource limitations [see, e.g. 4, 5, 6].
@@
Not all components in transportation networks are equally important because disruptions in some parts of the network cause its performance to degrade more than those in other parts. Moreover, because impacts on network performance depend on the states of all network components, there is a need to evaluate the impacts of combinations of disruptions (e.g., a component may not be so important on its own, but its disruption simultaneously with another component may have drastic consequences for the performance of the network). Furthermore, it is pertinent to examine how the expected impacts of disruptions depend on alternative assumptions about the underlying probabilities of component failures, given that the information about these probabilities may be incomplete.
@@
Fortification actions reduce the probability of disruptions that can undermine the performance of the network. Because many fortification actions are typically implemented jointly, we analyze portfolios consisting of many actions. These analyses support the cost-efficient allocation of resources to those portfolios of fortification actions, which, taken together, provide the highest expected performance relative to the costs of implementing these actions [7]. Furthermore, such analyses can be produced for a range of budget levels for aggregate implementation costs, for example, to prepare plans for the cost-efficient allocation of additional resources or, in case of budget cuts, to identify which one of the previously planned fortification actions can be left out so that the reduction on expected network performance will remain as small as possible.
@@
In this paper, we develop an approach based on multicriteria portfolio decision analysis in order (i) to assess and aggregate multiple objectives that reflect the services provided by infrastructure networks and (ii) to guide the allocation of resources to actions that are cost-efficient in mitigating disruptions in different parts of the network. Specifically, we represent the network by nodes (components subject to failure) and edges (connections between these components), and our formulation can be used to identify which nodes (or combinations of nodes) are most important to network performance. Moreover, it helps identify which portfolios of fortification actions are cost-efficient in ensuring the desired level of network performance. Examples of such fortification actions include fortifying selected nodes against disruptions.
@@
The remainder of this paper is organized as follows. Section 2 discusses earlier approaches to analyzing disruptions and their impacts on the performance of transportation networks. Section 3 develops an approach that combines probabilistic risk assessment with multicriteria portfolio decision analysis to quantify the expected performance of the network with regard to several objectives and, in addition, guides the allocation of resources to those fortification actions that fortify the network cost-efficiently. Section 4 presents a case study on the fortification of switches to improve the reliability of connections at a representative railway station in Finland. Section 5 discusses the numerical results, and Section 6 outlines future research areas.
@@ 2 Background
Early studies of railway networks focused mainly on their topological configuration due to limited data and computational power [8]. Later, there was a proliferation of specific models developed to reduce travel times [9], design schedules [10], plan new lines [11], reduce operating costs [12], and ensure track functionality, among other objectives. In many countries, infrastructure decisions (e.g., ensuring that the railway network is safe and functional) and operational decisions (e.g., ensuring that train timetables are being maintained) are made by entirely different entities. This is reflected in the fact that infrastructure and operational research have historically been somewhat decoupled. Within these streams of research, we focus on analyzing the network as a critical infrastructure in providing services.
@@
Most railway system analysis relies on network theory to identify critical components, evaluate network performance, and develop strategies to reinforce or disrupt them [see, e.g., 13, 14, 15, 16]. Latora and Marchiori [17] present a method to find the critical components of an infrastructure network represented by nodes and edges. They also analyze how improvements, such as adding edges between nodes, enhance network performance. In their model, performance is measured using a topological metric, i.e., a metric that relies purely on how nodes are positioned and connected in the network, to quantify how efficiently information can be exchanged over the network. Ip and Wang [18] propose a methodology to evaluate the resilience of railway networks based solely on topological metrics, such as the number of independent paths. They also propose an algorithm to fortify the network to increase its resilience. Fecarotti et al. [19] propose a non-linear integer programming model that considers topological metrics to select maintenance strategies.
@@
Although several studies have used topological metrics to assess network performance, few have evaluated the quality of these assessments [20]. In this regard, a general drawback of topological metrics is that these metrics are not necessarily closely related to network performance objectives. For example, Hao et al. [21] propose a multi-objective optimization approach to identify critical components in a network, accounting for their interactions with other systems. One of their key findings is that the criticality of the nodes is not necessarily related to their topological importance. LaRocca et al. [22] conclude that many topological metrics are of limited value in analyzing the robustness of power systems in different disruption scenarios. Moreover, Alderson et al. [23] show that the criticality of a given component can depend more on the set of disrupted components than on the component itself.
@@
Another challenge in evaluating the performance of networks under uncertainty is the identification of possible hazards and their impacts. In some cases, it is unclear what hazards can affect the network [24] or to what extent the infrastructure will withstand expected / unexpected hazards [25]. Zhang et al. [26] present a summary of recent studies to quantify the loss of functionality of railway systems due to various hazards, such as extreme weather or seismic events. They propose a framework for combined hazards to simulate functionality loss in a coupled railway and airline system. One of the key aspects in quantifying this loss is to account for the propagation of the disruptions. In general, the assumption of independent failure rates can lead to overestimating the reliability of the network, leading to poor design and unacceptable performance. Nazarizadeh et al. [27] propose a model which includes common cause failures and interactive failures. They apply the model to the Iranian railway system and show that it can provide better reliability estimates than other methodologies.
Several authors proceed by elaborating scenarios characterizing the realization of hazards and estimating their probabilities. For example, Joshi et al. [28] and Yang et al. [29] consider scenarios of rainfall and tornadoes to assess the risks affecting the railway systems in India and China, respectively. Turoff et al. [30] propose a collaborative dynamic scenario model based on expert judgments to estimate the cascading effects of critical infrastructure interactions.
Zio [25] and Sedghi et al. [31] call for the development of frameworks that help railway infrastructure managers understand and quantify the complexity of the network and, by doing so, help them prepare for hazards to ensure acceptable network performance. In this context, we develop a general assessment approach and accompanying search algorithms that help assess the importance of components in a rail network, considering disruption probabilities, multiple objectives associated with the services provided by the network, and possible minimum performance requirements on these objectives. This assessment of the components’ importance is further used to provide fortification recommendations.
3 Proposed Approach for Fortifying Networks
@@ 3.1 Network Representation
Let $G(V;E)$ denote a network consisting of a set of nodes $V=\{1,\dots,n\}$ and a set of undirected edges $E\subseteq\{(i,i^{\prime})\mid i,i^{\prime}\in V\}$ between the nodes. A path is a sequence of nodes and edges that connect two nodes in the network. The state of a node is operational or disrupted. If a node is disrupted, none of the paths containing it can be traversed. Thus, if the nodes in $D\subseteq V$ are disrupted, the remaining network is $G(V^{D};E^{D})$, where $V^{D}=V\setminus D$ and $E^{D}=\{(i,i^{\prime})\in E\mid i,i^{\prime}\in V^{D}\}\subseteq E$.
The state $x_{k}$ of node $k$ is modeled as a realization of a binary random variable $X_{k}$ such that $x_{k}=0$ if node $k\in V$ is disrupted and $x_{k}=1$ if it is operational. The state of the network $x=(x_{1},\dots,x_{n})\in\mathcal{X}=\{0,1\}^{n}$ is the joint realization of the random variables that represent the states of nodes. Thus, there are $2^{n}$ network states. We assume that the disruption events at the nodes occur independently and that $p_{k}=\mathbb{P}[X_{k}=0]$ is the probability that node $k$ is disrupted. Due to this independence assumption, the probability distribution over network states is fully characterized by the vector $p=(p_{1},\ldots,p_{n})$.
@@ 3.2 Assessing Network Performance
Critical infrastructure networks can enable multiple services to their users, so their performance is generally measured by considering several objectives. For example, the performance of a transportation network can be assessed with regard to its ability to provide connections between relevant origin-destination pairs. The objectives can also be assessed in terms of topological metrics such as the average of the shortest distances between the nodes of the network [see, e.g., 32, 33].
@@
The performance of the network depends on its state because if some nodes are disrupted, the network may not be able to provide its intended services. The attainment of $m$ multiple objectives can be measured by corresponding criteria and, more technically, by utility functions $u_{j}(\cdot):\mathcal{X}\mapsto[0,1]$ such that $u_{j}(x)$ represents the utility that is associated with the performance on criterion $j=1,\dots,m$ when the network is in state $x\in\mathcal{X}$. We assume that the network is coherent in that these utilities cannot increase with the disruption of a node. The utilities are normalized so that the minimum $u_{j}(x)=0$ is associated with the states that give the worst performance on criterion $j$ (i.e., all nodes are disrupted $x^{\circ}_{k}=0,k=1,\dots,n$). The maximum $u_{j}(x)=1$ is attained when the network performs at its best on the $j$-th criterion (i.e., all nodes are operational $x^{*}_{k}=1,k=1,\dots,n$).
The utility functions $u_{j}(\cdot)$ can be aggregated by employing the additive multicriteria utility function in (1), assuming that the criteria are mutually preferentially independent (i.e., preferences for a given criterion do not depend on those for any other criteria), and every criterion is additive independent (i.e., there are no preferences for how the realizations for a given criterion coincide with realizations with other criteria, provided that the probabilities of all realizations on the different criteria remain unchanged) [see 34]. Specifically, in the utility function
$u(x,w)=\sum_{j=1}^{m}w_{j}u_{j}(x)\in[0,1],$ (1)
the weight of the criterion $w_{j}\in[0,1],j=1,\dots,m$ reflects the relative overall utility increase gained as a result of the improvement on criterion $j$ when the state of the network changes from its worst state (all nodes are disrupted) to its best state (all nodes are operational). Following the usual convention, these weights can be normalized so that $\sum_{j=1}^{m}w_{j}=1$.
@@
However, obtaining a complete preference characterization through criteria weights can be challenging. For example, if the different criteria represent the planned volume of traffic between different origin-destination pairs in the network, it is possible that this volume is not fully known at the time of planning the fortification actions. In other cases, decision-makers (DM) may be reluctant to specify the importance of criteria, or there can be multiple DMs with different priorities. In response to this recognition, and also to support extensive sensitivity analyses, we characterize information about the relative importance of criteria through an information set $\mathcal{S}$ consisting of all the weights that are compatible with the elicited preference statements [35]. These statements can be elicited by asking the DM to express statements concerning the relative importance of criteria. Thus, for example, if criterion 1 is at least as important as criterion 2 but not more than two times more important, the constraints $w_{2}\leq w_{1}\leq 2w_{2}$ hold. Assuming that the preferences are elicited through statements that correspond to linear constraints, then the resulting weight set $\mathcal{S}$ is a subset of the non-informative weight set $\mathcal{S}^{0}$
@@
3.3 Fortifying the Network
The expected performance of the network can be improved by implementing fortification actions that, for example, lead to lower disruption probabilities at the nodes where they are implemented. Figure 1 illustrates the decision to implement an action to fortify a network node. Without this action, the probability of disruption of node $k$ is $p_{k}$. If the corresponding action is implemented, this probability is reduced to $p^{\prime}_{k}<p_{k}$.
Figure 1: Decision tree for implementing an action to fortify network node $k$.
Usually, there are many possible fortification actions for improving network performance. As a result, the probability of different levels of network performance depends on what is the portfolio (i.e., subset) of actions that are implemented. We assume that there are $h$ possible fortification actions, represented by the binary variables $q_{1},\ldots,q_{h}$ such that $q_{l}=1$ if action $l$ is implemented and $q_{l}=0$ if not, $l=1,\dots,h$. The portfolio of fortification actions is given by a vector $q=(q_{1},\dots,q_{h})\in\mathcal{Q}=\{0,1\}^{h}$. The total cost of a portfolio $q$ is $c_{tot}(q)\in\mathbb{R}_{+}$. This total cost $c_{tot}$ is taken to be the sum of the costs of the individual actions $c_{l}$ in the portfolio so that $c_{tot}(q)=\sum_{l=1}^{h}c_{l}q_{l}$. More complex cost structures can be easily introduced to account for cost synergies among actions.
A portfolio of fortification actions $q$ is feasible if (i) its total cost $c_{tot}(q)$ does not exceed the level of available resources $r$ (i.e., $c_{tot}(q)\leq r$) and (ii) satisfies relevant logical constraints (e.g., if actions $1$ and $2$ are mutually exclusive, allowing at most one of them to be implemented, the constraint $q_{1}+q_{2}\leq 1$ holds). The set of feasible portfolios is denoted by $\mathcal{Q}_{F}\subseteq\mathcal{Q}$. The aim is to determine those sets of feasible portfolios that satisfy the relevant constraints and contribute the most to the attainment of objectives, such as maximizing expected performance.